OpenStudy (anonymous):

Solve 4(x + 1) + 8 = 24. 3.75 -2 5 3

OpenStudy (solomonzelman):

4(x + 1) + 8 = 24 4(x + 1) = 16 x + 1 = 4 x=?

OpenStudy (anonymous):

subract from one on both sides and go from there

OpenStudy (anonymous):

its 3

OpenStudy (solomonzelman):

yes, 3

OpenStudy (anonymous):

Can you help me out with a few more?

OpenStudy (anonymous):

uh ummm sure

OpenStudy (anonymous):

Solve for x: 1 over 4 (2x − 14) = 4. 15 15 over 2 9 11

OpenStudy (anonymous):

im not sure on this one refer to your notes to help you

OpenStudy (anonymous):

(2x - 14) = 4 *** my bad

OpenStudy (anonymous):

add 14 on both sides

OpenStudy (anonymous):

it cancels the 14 and leavse it like 2x=18 and you get 9

OpenStudy (anonymous):

Solve for x: the quantity of x plus 4 all over 2 = 7. x = 3 x = 5 x = 6 x = 10

OpenStudy (anonymous):

is this one where you rewrite and then solve